NOAA SATOPS Morning Report: Monday, January 07, 2013


SATOPS Morning Report: Monday, January 07, 2013

Customers or users who want to call to ask about problems or report outages can use the following contacts:
- For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444
- To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

DOY 007

GOES
All GOES operations were nominal over the past 72 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
N19 DOY005 08:22z GAC3A was not transferred to Eumetsat due to a network problem. Under investigation.

All POES operations were nominal over the past 72 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
OPS54 DOY006 06:50Z CAT III data loss of 5322 s/c seconds. Under investigation.

OPS51 DOY007 02:33z data loss of 6057 SC seconds. Under investigation.

All other DMSP operations were nominal over the past 72 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All JASON-2 operations were nominal over the past 72 hours.
No anomalies affecting product delivery or s/c health & safety.

ESPC Reported data outage and delays
DOY005 ESPC did not receive N19 HPRT Rev. 20152 from WCDA due to a network problem.

DOY007 ESPC did not receive F-17 rev. 31861.

Questions or comments to Ron.Mahmot@noaa.gov

Please follow SpaceRef on Twitter and Like us on Facebook.